# Unveiling the Capabilities of Large Language Models in Detecting Offensive Language with Annotation Disagreement

### v1

@@ 4.2 Impact of Disagreement Samples on Instruction Fine-tuning
We select LLaMa3-7B as the backbone model and evaluate its performance after instruction fine-tuning. Specifically, following the settings of Leonardelli et al. (2021), we randomly sample an equal number of instances with different agreement levels from the training set of the MD-Agreement dataset for fine-tuning. We extract 1,800 samples each from the $A^{++}$, $A^{+}$, and $A^{0}$ categories, with this quantity based on the $A^{0}$ samples, as they are the least represented in the dataset. The instruction template remains consistent with that used in the zero-shot setting (see Appendix A.3). We also examine the impact of both single-category agreement samples and combinations of varying agreement degrees, adopting the same experimental markers as in Section 4.1. The results of the MD-Agreement test set are presented in Table 5. Based on these results, we can observe that:
(1) When fine-tuned with a single category of annotation agreement samples, LLaMA3-8B performs best with medium agreement $A^{+}$ samples, achieving strong classification accuracy and alignment with human annotations. In contrast, the effects of high-agreement $A^{++}$ samples and low-agreement $A^{0}$ samples explicitly differ: fine-tuning with $A^{++}$ enhances the model’s classification ability, while incorporating $A^{0}$ samples improves the alignment degree between LLM confidence and human annotation agreement, resulting in the lowest MSE. This phenomenon highlights the "double-edged sword” nature of disagreement samples in model detection performance. On the one hand, learning from these samples enables the model to capture the ambiguous boundary between offensive and non-offensive language while mitigating overconfidence in decision-making. On the other hand, relying solely on these challenging samples for fine-tuning may risk catastrophic forgetting, ultimately impairing overall performance. For this reason, $A^{+}$ samples provide the most optimal benefits, striking a well-balanced trade-off between effective learning and preventing model forgetting, enabling the model to enhance the ability without compromising overall performance due to excessive focus on complex samples.
(2) Fine-tuning with combinations of different agreement categories further improves the model’s detection performance compared to training with a single category of annotation agreement samples. Notably, training LLaMA3-8B with a mix of all three disagreement levels ($w/\;A^{++/+/0}$) achieves the best results, with overall performance comparable to GPT-4o in the few-shot learning (see Table 5). This underscores the advantage of training with a larger sample size, enabling a more comprehensive understanding and detection of offensive language. Additionally, among the combinations involving two categories, combining the two types of disagreement samples ($w/\;A^{+/0}$) shows explicitly better overall performance compared to the other two combinations, despite all of them using the same amount of instruction data. This further underscores the effectiveness of learning from disagreement samples and demonstrates that a well-balanced selection of different agreement levels during instruction fine-tuning can significantly benefit model performance.
We also replicate the instruction fine-tuning experiment with Qwen2.5-7B, using the same data for training and testing. The results confirm consistent conclusions, shown in Appendix B.4.
@@ 5 Related Work
Large Language Model. In recent years, large language models (LLMs) have rapidly emerged, showcasing extensive world knowledge and strong reasoning capabilities Kojima et al. (2022); Ouyang et al. (2022); OpenAI (2023). Many researchers have proposed diverse tasks to deeply analyze the relationship between the model’s outputs and human judgments Xu et al. (2024); Fan et al. (2024). In addition, the confidence of LLMs in their outputs has also attracted attention from researchers, which is often used to assess the reliability and robustness of the generated content Jiang et al. (2021). Various methods for estimating confidence have been proposed Zhang et al. (2020); Wang et al. (2023b); Tian et al. (2023); Lin et al. (2022). In this study, we employ the most straightforward approach, self-consistency, to estimate the model’s confidence.
Offensive Language Detection. Researchers have developed various methods for detecting offensive language Founta et al. (2018); Davidson et al. (2017); Mathew et al. (2021). As research advances, many studies argue that treating offensive language detection as a binary classification is an idealized assumption Basile et al. (2021); Basile (2020); Plank (2022), as annotation disagreement are inherent in datasets for such subjective task Pavlick and Kwiatkowski (2019); Uma et al. (2021b). Using majority voting for annotation agreement leads to information loss Davani et al. (2022), as these disagreements arise from the subtlety of the samples, not labeling errors Uma et al. (2022). Leonardelli et al. (2023) emphasizes that detection models should recognize this disagreement, rather than just improving classification performance.
@@
6 Conclusions
In this paper, we focus on the ability of LLMs to detect offensive language with annotation disagreement. We first evaluate the performance of multiple LLMs in a zero-shot scenario, examining both binary accuracy and the relationship between LLM confidence and annotation agreement. Compared to unanimous agreement samples, all LLMs exhibit explicitly lower detection accuracy for samples with annotation disagreement, failing to capture their ambiguous nature. Moreover, due to the inherent overconfidence, these models tend to make consistent decisions, resulting in a low degree of alignment with human annotators. These issues highlight the challenges posed by disagreement samples in offensive language detection and demonstrate that existing LLMs cannot fully replace humans in detection.
We further explore the impact of disagreement samples on LLM detection performance under few-shot learning and instruction tuning. Specifically, we analyze the effect of a single agreement degree and different combinations of agreement categories. Experimental results indicate that learning from disagreement samples notably improves detection accuracy and alignment degree with human annotations. This suggests that disagreement samples help the model learn the nuanced boundary between offensive and non-offensive language as defined by human society. Our study underscores the importance of disagreement samples in advancing offensive language detection with LLMs. We hope our findings and insights will contribute to the further development of related research.
@@ A.4 Other Experimental Settings
We access closed-source LLMs via their official APIs and deploy open-source LLMs with parameters downloaded from Hugging Face  [^2] https://huggingface.co/. To ensure a fair comparison, we use model versions released around the same time, as detailed in Table A4. Except for the temperature coefficient, other hyperparameters, such as top-p and top-k, are set to their default values for each model. For instruction fine-tuning, we adopt the efficient Qlora fine-tuning method. The learning rate is set to 2e-4, with a per-device batch size of 36. We train the model for 15 epochs using the AdamW optimizer, applying an early stopping mechanism. We reserve the parameters of best-performing models based on the development set and evaluate their performance on the test set. The models are trained on two NVIDIA H100 80GB GPUs. All the few-shot learning and instruction fine-tuning experiments are repeated five times with different random seeds to minimize error, and the average results are reported.
@@ A.5 Handling of Refusal Behavior
Handling offensive language can trigger the refusal behavior of LLMs, as they are designed with ethical and safety considerations Kumar et al. (2024). Nevertheless, in our experiments, refusal occurred only in the zero-shot evaluation setting, where Claude-3.5, with a temperature coefficient set to 1, failed to generate responses for 23 samples. When the experiment was repeated with the same settings, the model successfully provided predictions for these samples. This phenomenon also highlights the model’s sensitivity to offensive language and the non-deterministic nature of its outputs.

### v2

@@ 4 RQ2: Impact of Disagreement Samples on LLM Learning
4.1 Impact of Disagreement Samples on Few-Shot Learning
We evaluate the effect of disagreement samples on GPT-4o’s binary classification accuracy and its confidence alignment with human annotations during few-shot learning.
Few-Shot Learning Setup.
We evaluate both single-category agreement samples and combinations of agreement levels in few-shot learning, following Leonardelli et al. (2021). We first construct prompts using positive and negative sample pairs randomly drawn from the MD-Agreement training set, with each prompt including pairs corresponding to the respective agreement level. For example, the simplest setup $w/\;A^{++}$ consists of only unanimous agreement ($A^{++}$) samples, containing one offensive and one non-offensive example. Furthermore, we examine mixed setups with different agreement configurations, consisting of sample pairs from their respective single categories for reliable evaluation. For instance, $w/\;A^{++/0}$ and $w/\;A^{++/+}$ combine unanimous agreement samples with one level of disagreement, respectively. Additionally, we assess a broader configuration, $w/\;A^{++/+/0}$, which includes samples from all three agreement levels. The template details are provided in Appendix A.3.
We evaluate model performance on the MD-Agreement test set, analyzing both overall results and performance across different agreement levels. Table 6 summarizes the key findings.
(1) Few-shot learning improves classification accuracy but may increase overconfidence in ambiguous samples.
Few-shot learning enhances classification accuracy, particularly in the medium agreement subset ($A^{+}$), where accuracy increases by an average of 3.87%. However, for detection of low-agreement samples ($A^{0}$), few-shot learning increases the MSE, suggesting that models become overconfident and misaligned with ambiguous human annotations. This occurs because few-shot learning reinforces model consistency, making it less adaptable to subjective disagreements.
(2) Learning from disagreement samples improves model generalization.
Using disagreement samples ($A^{+}$ and $A^{0}$) in few-shot learning leads to greater performance improvements across all evaluation metrics compared to using only unanimous agreement samples ($A^{++}$). Disagreement samples often capture borderline or ambiguous cases, which challenge the model to refine its decision boundaries. Learning from these samples enhances the model’s ability to differentiate nuanced offensive language from non-offensive content.
(3) Combining different agreement levels enhances performance, but excessive variation reduces accuracy.
Incorporating both unanimous agreement samples and disagreement samples (e.g., $w/\;A^{++/0}$ or $w/\;A^{++/+}$) improves model performance compared to using only disagreement samples ($w/\;A^{+/0}$). However, including too many agreement categories ($w/\;A^{++/+/0}$) does not further enhance accuracy and may even decrease performance. The increased variation makes it harder for the model to establish clear decision boundaries, potentially leading to inconsistent classifications.
These results indicate that strategically balancing agreement levels is critical in few-shot learning. A well-chosen mix of clear and ambiguous cases helps the model generalize effectively, whereas excessive variation may introduce confusion and decrease performance.
To verify the robustness of our findings, we replicate the experiment using the open-source LLM Qwen2.5-72B. The results align closely with those of GPT-4o, suggesting that these insights generalize across different LLM architectures. Detailed results are provided in Appendix B.3.
@@ 4.2 Impact of Disagreement Samples on Instruction Fine-tuning
We analyze how instruction fine-tuning with different annotation agreement levels affects model performance, using LLaMa3-7B as the backbone.
Instruction Fine-tuning Setup.
We fine-tune an equal number of instances from each agreement level in the MD-Agreement dataset. Specifically, we extract 1,800 samples each from $A^{++}$, $A^{+}$, and $A^{0}$, based on the least-represented $A^{0}$ category. The instruction template remains consistent with that used in the zero-shot setting (see Appendix A.3). We also evaluate combinations of multiple agreement levels, using the same experimental markers as in Section 4.1. Table 5 presents the results, leading to the following conclusions:
(1) Medium-agreement ($A^{+}$) samples yield the best balance in fine-tuning.
Fine-tuning with high-agreement ($A^{++}$) samples improves classification accuracy, while low-agreement ($A^{0}$) samples enhance confidence alignment with human annotations, reducing MSE. However, exclusive reliance on $A^{0}$ samples may lead to catastrophic forgetting, where the model becomes overly attuned to ambiguous cases at the cost of general classification accuracy. $A^{+}$ samples offer the best trade-off, allowing the model to capture nuanced decision boundaries while maintaining robust performance.
(2) Combining multiple agreement levels further enhances performance.
Fine-tuning with all three agreement levels ($w/\;A^{++/+/0}$) achieves the best overall results, yielding performance comparable to GPT-4o in few-shot learning (see Table 5). Among two-category combinations, mixing disagreement samples ($w/\;A^{+/0}$) provides the most improvement, reinforcing the importance of disagreement-aware learning.
These results confirm that strategically selecting disagreement samples is essential for instruction fine-tuning. A well-balanced combination enhances both classification performance and confidence calibration, ensuring better alignment with human judgments.
We replicate the instruction fine-tuning experiment with Qwen2.5-7B using the same training and test data. The results closely align with those of LLaMa3-7B, confirming that these insights generalize across different model architectures. See Appendix B.4 for details.
@@ 5 Related Work
Large Language Model.
In recent years, large language models (LLMs) have rapidly emerged, showcasing extensive world knowledge and strong reasoning capabilities Kojima et al. (2022); Ouyang et al. (2022); OpenAI (2023). Many researchers have proposed diverse tasks to deeply analyze the relationship between the model’s outputs and human judgments Xu et al. (2024); Fan et al. (2024). In addition, the confidence of LLMs in their outputs has also attracted attention from researchers, which is often used to assess the reliability and robustness of the generated content Jiang et al. (2021). Various methods for estimating confidence have been proposed Zhang et al. (2020); Wang et al. (2023b); Tian et al. (2023); Lin et al. (2022). In this study, we employ the most straightforward approach, self-consistency, to estimate the model’s confidence.
Offensive Language Detection.
Researchers have developed various methods for detecting offensive language Founta et al. (2018); Davidson et al. (2017); Mathew et al. (2021). As research advances, many studies argue that treating offensive language detection as a binary classification is an idealized assumption Basile et al. (2021); Basile (2020); Plank (2022), as annotation disagreement are inherent in datasets for such subjective task Pavlick and Kwiatkowski (2019); Uma et al. (2021b). Using majority voting for annotation agreement leads to information loss Davani et al. (2022), as these disagreements arise from the subtlety of the samples, not labeling errors Uma et al. (2022). Leonardelli et al. (2023) emphasizes that detection models should recognize this disagreement, rather than just improving classification performance.
@@
6 Conclusion
This study examines how LLMs handle annotation disagreement in offensive language detection, a critical challenge in real-world moderation. We evaluate multiple LLMs in a zero-shot setting and find that while they perform well on unanimously agreed-upon samples, their accuracy drops significantly for disagreement cases. Moreover, their overconfidence leads to rigid predictions, misaligning them with human annotations.
To address this, we investigate the impact of disagreement samples in few-shot learning and instruction fine-tuning. Our results show that incorporating these samples improves detection accuracy and human alignment, enabling LLMs to better capture the subjective nature of offensive language. We further find that balancing agreement levels in training data prevents overfitting to ambiguous cases, ensuring model robustness.
Key findings of this work include: (1) a systematic evaluation of LLMs on annotation disagreement, (2) insights into how disagreement samples improve learning, and (3) guidelines for leveraging disagreement-aware training strategies. These results emphasize the need for model calibration techniques to mitigate overconfidence and for training strategies that incorporate disagreement to improve generalization. Future research should explore dynamic fine-tuning approaches and confidence-aware moderation systems to bridge the gap between LLM decisions and human subjectivity.
@@ A.4 Other Experimental Settings
We access closed-source LLMs via their official APIs and deploy open-source LLMs with parameters downloaded from Hugging Face. To ensure a fair comparison, we use model versions released around the same time, as detailed in Table A4. Since GPT-o1 only has a default temperature of 1 and does not allow adjustments, we present its binary performance in this setting. Except for the temperature coefficient, other hyperparameters, such as top-p and top-k, are set to their default values for each model. For instruction fine-tuning, we adopt the efficient Qlora fine-tuning method. The learning rate is set to 2e-4, with a per-device batch size of 36. We train the model for 15 epochs using the AdamW optimizer, applying an early stopping mechanism. We reserve the parameters of best-performing models based on the development set and evaluate their performance on the test set. The models are trained on two NVIDIA H100 80GB GPUs. All the few-shot learning and instruction fine-tuning experiments are repeated five times with different random seeds to minimize error, and the average results are reported.
@@ A.5 Handling of Refusal Behavior
Handling offensive language can trigger the refusal behavior of LLMs, as they are designed with ethical and safety considerations Kumar et al. (2024). Nevertheless, in our experiments, refusal occurred only in the zero-shot evaluation setting, where Claude-3.5, with a temperature coefficient set to 1, failed to generate responses for 23 samples. When the experiment was repeated with the same settings, the model successfully provided predictions for these samples. This phenomenon also highlights the model’s sensitivity to offensive language.